Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night report 2017-11-25

Observers: Z.Kostrzewa-Rutkowska, K.Lopez
TAT: J. Chen, M. Fraser, J. Sollerman
Support team: P. Jonker, G. Cannizzaro
Instruments: SOFI, EFOSC2

Weather conditions: 
evening UT 23:45: thin clouds, humidity 35%
morning UT 8:30: clear, but clouds above the horizon, humidity 25%

Calibrations:
EFOSC2 bias, arcs daytime, arcs morning

LOG:

UT      Target           Type       Filter/Grism ExpTime[s]  Slit   Seeing  Airmass    Comments
--------SOFI---------
0:02    Feige110         STD        BGRG                            1.38    1.10
0:15    AT2017gww        Followup   JHK                             1.35    1.25  
0:47    AT2017hcc        Followup   JHK                             0.95    1.05       Issues with image quality - aborted
1:02    AT2017hcc        Followup   JHK                             1.10    1.06       Restarted and aborted
1:05    AT2017hcc        Followup   JHK                             0.91    1.06       Restarted
1:10    Hip80            Telluric   BGRG                            0.91    1.07
1:15    AT2017hcc        Followup   BGRG                            0.94    1.08
2:23    AT2016cvk        Followup   JHK                             0.88    1.44
3:44    AT2016jbu        Followup   JHK                             0.94    1.66 
5:00                                                                                   SOFI --> EFOSC2
-------EFOSC2--------
5:20    L745a            STD        gr11/gr13/gr16          1.0/1.5 0.81    1.29
5:45    MASTERJ034412..  Followup   gr13                    1.0     0.98    1.09
6:25    AT2016jbu        Followup   BVri                            1.02    1.35
6:40    AT2016jbu        Followup   gr16                    1.0     1.02    1.32
7:28    AT2016jbu        Followup   gr11                    1.0     0.87    1.31
8:10    L745a            STD        gr11/gr13/gr16          1.0/1.5 0.93    1.02         
Comments